What is the interval in which f(x)=(x^2-3x +2) / (x^2 + x -6) lies ?

Grade:11

1 Answers

sahil
6 years ago
Sol:If you will look it can be factorized, here f(x)=x^2-3x+2/x^2+x-6 or (x-2)(x-1)/(x-2)(x+3) ;cancel (x-2) to get f(x)=(x-1)/(x+3) ;Now i think you can go ahead with graph and find range.Hope it helps please approve the answer .